# Effect of interatomic repulsion and quasi-degenerate states on a Kitaev-transmon qubit based on double quantum dots

### v3

@@ Abstract
We investigate the effect of interatomic Coulomb repulsion $V$ and particular states disregarded previously on the Kitaev-transmon system proposed by Pino et al. [43] which consists of a Josephson junction between two double quantum dots (DQDs) modeled by the spinless Kitaev Hamiltonian. For an isolated DQD, we demonstrate that a “sweet spot” hosting “poor man’s Majorana” states persist in the presence of $V$, provided that system parameters are appropriately tuned. For the full system, we demonstrate that at the sweet spots of both DQDs, all eigenstates are doubly degenerate. This degeneracy arises from the existence of an operator that maps between two decoupled Hilbert subspaces. Away from the sweet spots, the microwave spectrum becomes sensitive to the choice of initial state. In our study, we consider transitions from the ground state (which depending on the flux alternates between the above mentioned subspaces) to all possible excited states. This scenario corresponds to a system initially in thermal equilibrium at low temperature.
I Introduction
@@
In recent years, topological superconducting systems have recently attracted great interest due the presence of Majorana zero modes (MZMs),[1, 2, 3, 4, 7, 5, 6] which are candidates to store and manipulate quantum information due to their non-Abelian exchange statistics [8]. Many studies were based on the Kitaev model for $p$-wave superconductors [9]. However, due to difficulties in realizing such a model experimentally, other systems were studied, such as nanowires with strong spin-orbit interaction with induced superconductivity by proximity to an $s$-wave superconductor [10, 11, 12, 13, 14, 15, 16],. After the first theoretical proposals [10, 11], MZMs were reported experimentally [12, 13, 14, 15]. In recent years, the phase diagram of the model [17] has been studied, along with its derived properties and implications. A few examples are in Refs. [18, 19, 20, 21, 22, 23, 24, 25].
@@
The system consisting of a topological superconducting wire and a quantum dot (QD) has also been studied. [15, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In particular Prada et al. proposed that a QD at the end of the nanowire may be used as a powerful spectroscopic tool to quantify the degree of Majorana nonlocality through a local transport measurement [23], and this proposal has been confirmed experimentally [28].
@@
Artificial Kitaev chains based on QDs have been proposed to circumvent disorder and other issues [32, 33, 34, 35, 36]. However, in these systems the topological protection is lost and the MZMs are obtained by fine tuning parameters such as the crossed Andreev reflection (CAR) and single-electron elastic cotunneling (ECT). For these reason these modes are called “poor man’s Majoranas” (PMMs) [32]. The point in the space of parameters at which these PMMs exists is called “sweet spot” [35]. The control of these parameters has been demonstrated experimentally in systems of two [37, 38] and three [39, 40] QDs. As a consequence, systems with PMMs in double QDs (DQDs) have been theoretically studied recently [41, 42, 43, 44, 45, 46].
@@
In particular, Pino et al. studied a qubit based on a Josephson junction between two DQDs described by the spinless Kitaev model (a minimal Kitaev-transmon qubit) [43]. They calculated the microwave spectrum as in similar systems studied previously [47, 48, 49, 50].
@@
The transmon qubits have received great attention in recent years [51, 52, 53, 54, 55, 56, 57, 58]. They present several advantages to alternative qubits, like reduced sensitivity to charge noise [57], long coherence times [51, 52], and strong coherent qubit-qubit coupling [56]. Some relevant experiments are in Refs. [51, 52, 53, 54, 55, 56]. A related model for a spinless Kitaev chain has been investigated in recent work [59], and the concept of Majorana-based transmons has been examined in earlier studies [60]. The expected key advantage of the Kitaev-transmon over the ordinary transmon is its inherent topological protection and robustness of the Majorana fermions. This design suppresses its sensitivity to dielectric loss, charge noise, and quasiparticles, which are expected to significantly extend the relaxation ($T_{1}$) and coherence ($T_{2}$) times.
@@
In Section II we study the conditions under which PMMs exist in an interacting DQD. In Section III we discuss the eigenstates and energies of a system of two DQDs connected by a Josephson junction. In Section IV we study the spectrum of the whole system, the Kitaev-transmon qubit, by numerically diagonalizing the Hamiltonian for some parameters of interest. Section V contains a summary.
II Double quantum dot with interdot repulsion
@@
$H_{\text{DQD}}=-\sum_{i}\mu_{i}c_{i}^{\dagger}c_{i}+(\Delta c_{1}c_{2}-tc_{1}^{\dagger}c_{2}+\text{H.c.})+Vc_{1}^{\dagger}c_{1}c_{2}^{\dagger}c_{2},$ (1)
@@
The parity of the number of particles is conserved and the Hamiltonian can be easily diagonalized in each subspace of well defined parity. In the subspace of even parity, the ground state and its energy are given by
$|g_{E}\rangle$ $=$ $u_{E}|0\rangle+v_{E}c_{1}^{\dagger}c_{2}^{\dagger}|0\rangle,\text{ }E_{E}=-\mu+\frac{V}{2}-r_{E},$
$u_{E}^{2}$ $=$ $\frac{1}{2}+\frac{V-2\mu}{4r_{E}},\text{ }v_{E}^{2}=1-u_{E}^{2},$
$\mu$ $=$ $\frac{\mu_{1}+\mu_{2}}{2},\text{ }r_{E}=\sqrt{\left(-\mu+\frac{V}{2}\right)^{2}+\Delta^{2}},$ (2)
with $u_{E},v_{E}>0$. For odd parity
$|g_{O}\rangle$ $=$ $u_{O}c_{1}^{\dagger}|0\rangle+v_{O}c_{2}^{\dagger}|0\rangle,\text{ }E_{O}=-\mu-r_{O},$
$u_{O}^{2}$ $=$ $\frac{1}{2}+\frac{\delta}{2r_{O}},\text{ }v_{O}^{2}=1-u_{O}^{2},$
$\delta$ $=$ $\frac{\mu_{1}-\mu_{2}}{2},\text{ }r_{O}=\sqrt{\delta^{2}+t^{2}},$ (3)
with $u_{O}>0$, $(v_{O})=sgn$$(t)$.
As discussed in Refs. [35, 45], a system hosting poor man’s Majoranas (PMMs) should satisfy four conditions, outlined below. To motivate these requirements for the unfamiliar reader, we briefly recall the defining properties of a genuine Majorana fermion. A Majorana operator is its own antiparticle and can be expressed as $\gamma=f+f^{\dagger}$, where $f$ is an ordinary fermion. Therefore it is chargeless. We assume that $\gamma$ is a one-particle eigenoperator of the Hamiltonian that describes the system, i.e. $\left[H,\gamma\right]=E_{\gamma}\gamma$. Taking the Hermitian conjugate of this relation one obtains the same equation with the opposite sign of $E_{\gamma}$ implying $E_{\gamma}=0$. Therefore, if the ground state for an even number of particles is denoted by $|E\rangle$, then the state $|O\rangle=\gamma|E\rangle$ is a state with an odd number of particles and the same energy. Note that since $\gamma^{2}=1$, $|E\rangle=\gamma|O\rangle$. For PMMs, one seeks to replicate these key properties of Majoranas, with two additional requirements: perfect localization at a single quantum dot (the third condition, essential for braiding operations), and the presence of a finite excitation gap, ensuring robustness of the system against disorder and small perturbations (fourth condition). If $\gamma$ is localized at one site $j$, it commutes with the occupancy operator $c_{k}^{\dagger}c_{k}$ for $k\neq j$. Therefore, the charge under addition of $\gamma$ is conserved at all sites rather than merely conservation of the total charge.
@@
$|\langle g_{E}|\tilde{\gamma}_{j}|g_{O}\rangle|=1,\text{ }|\langle g_{E}|\tilde{\gamma}_{k}|g_{O}\rangle|=0$ (5)
@@
$\Delta Q_{1}$ $=$ $\frac{2\mu-V}{4r_{E}}-\frac{\delta}{2r_{O}}=0,$
$\Delta Q_{2}$ $=$ $\frac{2\mu-V}{4r_{E}}+\frac{\delta}{2r_{O}}=0.$ (6)
This implies $\delta=0$, $\mu=V/2$, or $\mu_{1}=\mu_{2}=V/2$. These requirements for the sweet spot can be obtained tuning the gate voltages. The remaining requirement comes from the first condition, which is satisfied if $\Delta=|t|+V/2$. Recent results arrived to the same conclusion [70].
@@
Summarizing this Section, in presence of interdot interaction $V$, a sweet spot similar to the case $V=0$ can be reached, leading to well defined PMMs. This is defined by $\mu_{1}=\mu_{2}=V/2$ and $\Delta=|t|+V/2.$
III Two DQDs connected by a Josephson junction
@@
$H_{\text{4D}}=H_{\text{DQD}}^{L}+H_{\text{DQD}}^{R}+H_{J}.$ (8)
A scheme of the system is represented in Fig. 1. For simplicity we assume equal parameters for both dots, and that they are at the sweet spot except that $\Delta$ can be different from $t+V/2$. We also take $t>0$ (the sign of $t$ can be changed by the substitution $c_{L1}\rightarrow-c_{L1}$, $c_{R2}\rightarrow-c_{R2}$). In this case, the ground state for $t_{J}=0$ is either $|g_{E}\rangle$ for both dots (a state which we denote by $|ee\rangle$) if $\Delta>t+V/2$ or $|g_{O}\rangle$ for both dots ($|oo\rangle$) if $\Delta<t+V/2$. Both states have an even total number of particles. In more general situations, the ground state can have a total odd number of particles.
The term $H_{J}$ mixes the subspace of $|g_{E}\rangle$ for both dots with that of $|g_{O}\rangle$ for both dots. Calculating the matrix element between these two states leads to the matrix
$H_{\text{4D}}(\phi)=\left(\begin{array}[]{cc}E_{ee}&-\frac{t_{J}}{2}\cos{\left(\frac{\phi}{2}\right)}\\ -\frac{t_{J}}{2}\cos{\left(\frac{\phi}{2}\right)}&E_{oo}\end{array}\right),$ (9)
@@
IV The Kitaev-transmon qubit
@@
In this Section we add to the previous setup a superconducting wire containing another Josephson junction with characteristic energy $E_{J}$ and charging energy $E_{C}$, connecting the superconducting parts of the double quantum dots closing the circuit. A scheme of the system is shown in Fig. 2. The total Hamiltonian is [43]
$H$ $=$ $H_{\text{4D}}(\hat{\phi}-\phi_{ext})+H_{C}+H_{J}^{\prime},$
@@
In contrast to the previous section, here the phase $\hat{\phi}$ is treated as an operator conjugate to the operator $\hat{m}$, which is defined such that the number of Cooper pairs on the left and right sides of the junction are $m_{L}-\hat{m}$ and $m_{R}+\hat{m}$. In this framework, $\hat{m}$ increases or decreases by 1 when a Cooper pair tunnels across the junction. The state $\hat{m}=0$ corresponds to charge neutrality on both sides. Consequently, $\hat{m}=-i\partial/\partial\hat{\phi}$ [80, 81]. $m_{g}$ is controlled by an applied gate voltage. The phase $\phi_{ext}$ is related to the external magnetic flux $\Phi$ through the circuit by $\Phi=\Phi_{0}\phi_{ext}/(2\pi)$, where $\Phi_{0}=h/(2e)$ is the flux quantum.
For convenience we will use the operator $\hat{n}=2\hat{m}$, and define $\hat{\theta}=\hat{\phi}/2$ (so that $\hat{n}$ changes by 1 when a single particle crosses the junction). Note that the commutation relation $\hat{\phi},\hat{m}]=[\hat{\theta},\hat{n}]=i$ retains the same form with the new variables. The low-energy states of the system now have the form $|een\rangle$ ($|oon\rangle)$ if each DQD has an even (odd) number of particles, for a total $n$ (eigenvalue of $\hat{n}$) between the number of particles of both islands. Now using $e^{ib\hat{\theta}}|n\rangle=|n+b\rangle$ (this can be demonstrated in a similar way as for the traditional translation operator in quantum mechanics), it is clear that in the new basis
$H_{J}^{\prime}=-\frac{E_{J}}{2}\sum_{n=-\infty}^{\infty}\sum_{p}|ppn+2\rangle\langle ppn|+\text{H.c.},$ (11)
@@
In a similar way, for the off-diagonal part of $H_{\text{4D}}(\hat{\phi}-\phi_{ext})$
$2\cos(\hat{\theta}-\theta_{ext})=\sum_{n=-\infty}^{\infty}[(\cos(\theta_{ext})-i\sin(\theta_{ext}))$
$\times(|oon+1\rangle\langle een|+|een+1\rangle\langle oon|)+\text{H.c.}]$ (12)
@@
We have numerically diagonalized $H$ truncating the basis to $|n_{i}|<16$, which is enough to obtain accurate results. We take parameters similar to Ref. [43]: $t_{J}=E_{J}=E_{c}=1$, $\mu_{1}=\mu_{2}=V/2$, $E_{O}=-t-V/2=E_{oo}/2=1$, and vary $E_{E}=-\Delta=E_{ee}/2$.
Figure 4: Energies as a function of the offset charge for $\phi_{ext}=0$ and $E_{ee}=E_{oo}/2$ (top) and $E_{ee}=3E_{oo}/2$ (bottom). Other parameters as in Fig. 3
@@
$\tau_{z}=\sum_{n}|{ee,n}\rangle\langle{ee,n}|-|{oo,n}\rangle\langle{oo,n}|$ (13)
@@
$S(\omega)=\sum_{l}=|\langle l|\hat{n}|g\rangle|^{2}\delta(\omega+E_{g}-E_{l}),$ (14)
@@
For parameters corresponding to the double degeneracy of all levels, $E_{oo}-E_{ee}=\phi_{ext}=0$, our results are similar to those presented in Fig. 3 (h) of Ref. [43]. Instead, in the general case, the crossing of levels mentioned above change the ground state and affect the spectrum. In Fig. 5 we show the frequencies and intensities of the microwave spectrum for the same parameters as in Fig. 4. While our results are qualitatively similar to those of Ref. [43], there are differences in the regions $m_{g}<0.25$ and $m_{g}>0.75$ for $E_{ee}=E_{oo}/2$, and in the interval $0.25<m_{g}<0.75$ for $E_{ee}=3E_{oo}/2$ because of the change in the ground state.
@@
V Summary
We have demonstrated that sweet spots with well defined PMMs persist in DQDs described by the Kitaev model in presence of interatomic Coulomb repulsion $V$. Achieving this requires shifting the chemical potential of both dots and increasing the CAR amplitude $\Delta$ (or decreasing the ECT amplitude if possible) by $V/2$. This might be difficult for large $V$ indicating that interatomic repulsion is detrimental for achieving the sweet spots but do not necessarily renders them impossible.
These results provide a simple extension of the results of Pino et al. [43] for the Kitaev transmon containing two Josephson junctions and a charge reservoir (see Fig. 2) by incorporating Coulomb repulsion through simple parameter renormalization.
Outside of the sweet spots, there is an influence of states previously disregarded because they do not affect the essential physics of the qubit, specifically, $|een\rangle$ with $n$ even and $|oon\rangle$ with $n$ odd. We have calculated the microwave spectrum for some selected cases which display the effect of these states. This spectrum might be used to detect the sweet spots.
At the sweet spots of both DQDs (assumed identical), all states are doubly degenerate.

### v4

@@ Abstract
We investigate the effect of interatomic Coulomb repulsion $V$ and particular states disregarded previously on the Kitaev-transmon system proposed by Pino et al. [43] which consists of a Josephson junction connecting two double quantum dots (DQDs) modeled by the spinless Kitaev Hamiltonian. For an isolated DQD, we demonstrate that a “sweet spot” hosting “poor man’s Majorana” states persist in the presence of $V$, provided that system parameters are appropriately tuned. For the full system, we demonstrate that, at the sweet spots of both DQDs, all eigenstates are doubly degenerate. This degeneracy arises from the existence of an operator that maps between two decoupled Hilbert subspaces. Away from the sweet spots, the microwave spectrum becomes sensitive to the choice of initial state of the system. In our study, we consider transitions from the ground state (which depending on the flux alternates between the above mentioned subspaces) to all possible excited states. This scenario corresponds to a system initially in thermal equilibrium at low temperature.
keywords:
Kitaev transmon, interatomic Coulomb repulsion, microwave spectrum
articletype: Paper
1 Introduction
@@
In recent years, topological superconducting systems have attracted great interest due to the presence of Majorana zero modes (MZMs) [1, 2, 3, 4, 7, 5, 6], which are candidates to store and manipulate quantum information due to their non-Abelian exchange statistics [8]. Many studies were based on the Kitaev model for $p$-wave superconductors [9]. However, due to difficulties in realizing such a model experimentally, other systems were studied, such as nanowires with strong spin-orbit interaction with induced superconductivity by proximity to an $s$-wave superconductor [10, 11, 12, 13, 14, 15, 16]. After the first theoretical proposals [10, 11], MZMs were reported experimentally [12, 13, 14, 15]. In recent years, the phase diagram of the model [17] has been studied, along with its derived properties and implications. A few examples are in Refs. [18, 19, 20, 21, 22, 23, 24, 25].
@@
The system consisting of a topological superconducting wire and a quantum dot (QD) has also been studied [15, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In particular Prada et al. proposed that a QD at the end of the nanowire may be used as a powerful spectroscopic tool to quantify the degree of Majorana nonlocality through a local transport measurement [23], and this proposal has been confirmed experimentally [28].
@@
Artificial Kitaev chains based on QDs have been proposed to circumvent disorder and other issues [32, 33, 34, 35, 36]. However, in these systems the topological protection is lost and the MZMs are obtained by fine tuning parameters such as the crossed Andreev reflection (CAR) and single-electron elastic cotunneling (ECT). For this reason these modes are called “poor man’s Majoranas” (PMMs) [32]. The point in the space of parameters at which these PMMs exist is called “sweet spot” [35]. The control of these parameters has been demonstrated experimentally in systems of two [37, 38] and three [39, 40] QDs. As a consequence, systems with PMMs in double QDs (DQDs) have been theoretically studied recently [41, 42, 43, 44, 45, 46].
@@
In particular, Pino et al. studied a qubit based on a Josephson junction connecting two DQDs described by the spinless Kitaev model (a minimal Kitaev-transmon qubit) [43]. They calculated the microwave spectrum as in similar systems studied previously [47, 48, 49, 50].
@@
The transmon qubits have received great attention in recent years [51, 52, 53, 54, 55, 56, 57, 58]. They present several advantages over other qubit architectures, like reduced sensitivity to charge noise [57], long coherence times [51, 52], and strong coherent qubit-qubit coupling [56]. Some relevant experiments are in Refs. [51, 52, 53, 54, 55, 56]. A related model for a spinless Kitaev chain has been investigated in recent work [59], and the concept of Majorana-based transmons has been examined in earlier studies [60]. The expected key advantage of the Kitaev-transmon over the ordinary transmon is its inherent topological protection and robustness of the Majorana fermions. This design suppresses its sensitivity to dielectric loss, charge noise, and quasiparticles, which are expected to significantly extend the relaxation ($T_{1}$) and coherence ($T_{2}$) times.
@@
In Section 2 we study the conditions under which PMMs exist in an interacting DQD. In Section 3 we discuss the eigenstates and energies of a system of two DQDs connected by a Josephson junction. In Section 4 we study the spectrum of the whole system, the Kitaev-transmon qubit, by numerically diagonalizing the Hamiltonian for some parameters of interest. Section 5 contains a summary.
2 Double quantum dot with interdot repulsion
@@
$H_{\mathrm{DQD}}=-\sum_{i}\mu_{i}c_{i}^{\dagger}c_{i}+(\Delta c_{1}c_{2}-tc_{1}^{\dagger}c_{2}+\mathrm{H.c.})+Vc_{1}^{\dagger}c_{1}c_{2}^{\dagger}c_{2},$ (1)
@@
The parity of the number of particles is conserved and the Hamiltonian can be easily diagonalized in each subspace of well defined parity. In the subspace of even parity, the ground state $|g_{E}\rangle$ and its energy $E_{E}$ are given by
$|g_{E}\rangle$ $=$ $u_{E}|0\rangle+v_{E}c_{1}^{\dagger}c_{2}^{\dagger}|0\rangle,\mathrm{}E_{E}=-\mu+\frac{V}{2}-r_{E},$
$u_{E}^{2}$ $=$ $\frac{1}{2}+\frac{V-2\mu}{4r_{E}},\mathrm{}v_{E}^{2}=1-u_{E}^{2},$
$\mu$ $=$ $\frac{\mu_{1}+\mu_{2}}{2},\mathrm{}r_{E}=\sqrt{\left(-\mu+\frac{V}{2}\right)^{2}+\Delta^{2}},$ (2)
where the coefficients of the ground state $u_{E},v_{E}>0$ and $\mu$ is the average of the chemical potentials. Similarly, for odd parity, changing the subscript $E$ (even) for $O$ (odd), the result is
$|g_{O}\rangle$ $=$ $u_{O}c_{1}^{\dagger}|0\rangle+v_{O}c_{2}^{\dagger}|0\rangle,\mathrm{}E_{O}=-\mu-r_{O},$
$u_{O}^{2}$ $=$ $\frac{1}{2}+\frac{\delta}{2r_{O}},\mathrm{}v_{O}^{2}=1-u_{O}^{2},$
$\delta$ $=$ $\frac{\mu_{1}-\mu_{2}}{2},\mathrm{}r_{O}=\sqrt{\delta^{2}+t^{2}},$ (3)
where $u_{O}>0$, $(v_{O})=sgn$$(t)$ and $2\delta$ is the difference between the chemical potentials.
As discussed in Refs. [35, 45], a system hosting poor man’s Majoranas (PMMs) should satisfy four conditions, outlined below. To motivate these requirements for the unfamiliar reader, we briefly recall the defining properties of a genuine Majorana fermion. A Majorana operator is its own antiparticle and can be expressed as $\gamma=f+f^{\dagger}$, where $f$ is an ordinary fermion. Therefore, it is chargeless. We assume that $\gamma$ is a one-particle eigenoperator of the Hamiltonian that describes the system, i.e. $\left[H,\gamma\right]=E_{\gamma}\gamma$. Taking the Hermitian conjugate of this relation one obtains the same equation with the opposite sign of $E_{\gamma}$ implying $E_{\gamma}=0$. Therefore, if the ground state for an even number of particles is denoted by $|E\rangle$, then the state $|O\rangle=\gamma|E\rangle$ is a state with an odd number of particles and the same energy. Note that since $\gamma^{2}=1$, $|E\rangle=\gamma|O\rangle$. For PMMs, one seeks to replicate these key properties of Majoranas, with two additional requirements: perfect localization at a single quantum dot (the third condition, essential for braiding operations), and the presence of a finite excitation gap, ensuring robustness of the system against disorder and small perturbations (fourth condition). If $\gamma$ is localized at one site $j$, it commutes with the occupancy operator $c_{k}^{\dagger}c_{k}$ for $k\neq j$. Therefore, the charge under addition of $\gamma$ is conserved at all sites rather than merely conservation of the total charge.
@@
$|\langle g_{E}|\tilde{\gamma}_{j}|g_{O}\rangle|=1,\>|\langle g_{E}|\tilde{\gamma}_{k}|g_{O}\rangle|=0$ (5)
@@
$\Delta Q_{1}=\frac{2\mu-V}{4r_{E}}-\frac{\delta}{2r_{O}}=0,\>\Delta Q_{2}=\frac{2\mu-V}{4r_{E}}+\frac{\delta}{2r_{O}}=0.$ (6)
This implies $\delta=0$, $\mu=V/2$, or $\mu_{1}=\mu_{2}=V/2$. These requirements for the sweet spot can be obtained by tuning the gate voltages. The remaining requirement comes from the first condition, which is satisfied if $\Delta=|t|+V/2$. The same equations were obtained in recent work; however, the four conditions required for the realization of PMMs were not analyzed in detail there [70].
@@
In summary, in the presence of an interdot interaction $V$, a sweet spot analogous to the case $V=0$ can still be achieved, leading to well-defined PMMs. This sweet spot is characterized by the conditions $\mu_{1}=\mu_{2}=V/2$ and $\Delta=|t|+V/2$. The first condition can be readily realized experimentally by tuning the gate voltages, whereas the second depends explicitly on the value of $V$ and may become difficult to satisfy for large $V$. Experimentally, the precise value of $V$ is generally unknown and may be challenging to estimate. From a theoretical perspective, however, interdot interactions are expected to play an important role in semiconductors with induced superconductivity and have therefore been included in several relevant studies [61, 62, 63, 64].
3 Two DQDs connected by a Josephson junction
@@
$H_{\mathrm{4D}}=H_{\mathrm{DQD}}^{L}+H_{\mathrm{DQD}}^{R}+H_{J}.$ (8)
A scheme of the system is represented in Fig. 1. For simplicity we assume equal parameters for both DQDs, and that they are at the sweet spot except that $\Delta$ can be different from $t+V/2$. We also take $t>0$ (the sign of $t$ can be changed by the substitution $c_{L1}\rightarrow-c_{L1}$, $c_{R2}\rightarrow-c_{R2}$). In this case, the ground state for $t_{J}=0$ is either $|g_{E}\rangle$ for both DQDs (a state which we denote by $|ee\rangle$) if $\Delta>t+V/2$ or $|g_{O}\rangle$ for both DQDs ($|oo\rangle$) if $\Delta<t+V/2$. Both states have an even total number of particles. In more general situations, the ground state can have a total odd number of particles.
The term $H_{J}$ mixes $|ee\rangle$ with $|oo\rangle$. Calculating the matrix element between these two states leads to the matrix
$H_{\mathrm{4D}}(\phi)=\left(\begin{array}[]{cc}E_{ee}&-\frac{t_{J}}{2}\cos{\left(\frac{\phi}{2}\right)}\\ -\frac{t_{J}}{2}\cos{\left(\frac{\phi}{2}\right)}&E_{oo}\end{array}\right),$ (9)
@@
4 The Kitaev-transmon qubit
@@
In this Section we add to the previous setup a superconducting wire containing another Josephson junction with characteristic energy $E_{J}$ and together with a capacitive shunt with charging energy $E_{C}$, connecting the superconducting parts of the double quantum dots closing the circuit. A scheme of the system is shown in Fig. 2. The total Hamiltonian is [43]
$H$ $=$ $H_{\mathrm{4D}}(\hat{\phi}-\phi_{ext})+H_{C}+H_{J}^{\prime},$
@@
In contrast to the previous section, here the phase $\hat{\phi}$ is treated as an operator conjugate to the operator $\hat{m}$, which is defined such that the number of Cooper pairs on the left and right sides of the junction are $m_{L}-\hat{m}$ and $m_{R}+\hat{m}$. In this framework, the expectation value of $\hat{m}$ increases or decreases by 1 when a Cooper pair tunnels across the junction. The eigenstate of $\hat{m}$ with eigenvalue zero corresponds to charge neutrality on both sides. Consequently, $\hat{m}=-i\partial/\partial\hat{\phi}$ [80, 81]. The quantity $m_{g}$ accounts for a charge imbalance controlled by an applied gate voltage. The phase $\phi_{ext}$ is related to the external magnetic flux $\Phi$ through the circuit by $\Phi=\Phi_{0}\phi_{ext}/(2\pi)$, where $\Phi_{0}=h/(2e)$ is the flux quantum.
For convenience we will use the operator $\hat{n}=2\hat{m}$, and define $\hat{\theta}=\hat{\phi}/2$ (so that $\hat{n}$ changes by 1 when a single particle crosses the junction). Note that the commutation relation $\hat{\phi},\hat{m}]=[\hat{\theta},\hat{n}]=i$ retains the same form with the new variables. The low-energy states of the system now have the form $|een\rangle$ ($|oon\rangle)$ if each DQD has an even (odd) number of particles, and $n$ is the eigenvalue of $\hat{n}$. Now using $e^{ib\hat{\theta}}|n\rangle=|n+b\rangle$ (this can be demonstrated in a similar way as for the traditional translation operator in quantum mechanics), it is clear that in the new basis
$H_{J}^{\prime}=-\frac{E_{J}}{2}\sum_{n=-\infty}^{\infty}\sum_{p}|ppn+2\rangle\langle ppn|+\mathrm{H.c.},$ (11)
@@
In a similar way, for the off-diagonal part of $H_{\mathrm{4D}}(\hat{\phi}-\phi_{ext})$
$2\cos(\hat{\theta}-\theta_{ext})=\sum_{n=-\infty}^{\infty}[(\cos(\theta_{ext})-i\sin(\theta_{ext}))(|oon+1\rangle\langle een|+|een+1\rangle\langle oon|)+\mathrm{H.c.}]$ (12)
@@
We have numerically diagonalized $H$ truncating the basis to $|n_{i}|<16$, which is enough to obtain accurate results. We take parameters similar to Ref. [43] but adapted to the fact that $V$ can be different from zero: $t_{J}=E_{J}=E_{c}=1$, $\mu_{1}=\mu_{2}=V/2$, $E_{O}=-t-V/2=E_{oo}/2=1$, and vary $E_{E}=-\Delta=E_{ee}/2$.
Figure 4: Energies as a function of the offset charge for $\phi_{ext}=0$ and $E_{ee}=E_{oo}/2$ (top) and $E_{ee}=3E_{oo}/2$ (bottom). Other parameters as in Fig. 3.
@@
$\tau_{z}=\sum_{n}(|{ee,n}\rangle\langle{ee,n}|-|{oo,n}\rangle\langle{oo,n}|)$ (13)
@@
$S(\omega)=\sum_{l}|\langle l|\hat{n}|g\rangle|^{2}\delta(\omega+E_{g}-E_{l}),$ (14)
@@
For parameters corresponding to the double degeneracy of all levels, $E_{oo}-E_{ee}=\phi_{ext}=0$, our results are similar to those presented in Fig. 3 (e) of Ref. [43]. Instead, in the general case, the crossing of levels mentioned above changes the ground state and affects the spectrum. In Fig. 5 we show the frequencies and intensities of the microwave spectrum for the same parameters as in Fig. 4. While our results are qualitatively similar to those of Ref. [43], there are differences in the regions $m_{g}<0.25$ and $m_{g}>0.75$ for $E_{ee}=E_{oo}/2$, and in the interval $0.25<m_{g}<0.75$ for $E_{ee}=3E_{oo}/2$ because of the change in the ground state.
@@
5 Summary
We have demonstrated that sweet spots with well defined PMMs, satisfying the four conditions for their existence, persist in DQDs described by the Kitaev model in presence of interatomic Coulomb repulsion $V$. Achieving this requires shifting the chemical potential of both dots and increasing the CAR amplitude $\Delta$ (or decreasing the ECT amplitude if possible) by $V/2$. This might be difficult for large $V$ indicating that interatomic repulsion is detrimental for achieving the sweet spots but does not necessarily render them impossible.
These results provide a simple extension of the results of Pino et al. [43] for the Kitaev transmon containing two DQDs, two Josephson junctions and a charge reservoir (see Fig. 2) by incorporating Coulomb repulsion through simple parameter renormalization (replacing $t$ by $t+V/2$ and setting $\mu_{1}=\mu_{2}=V/2$).
Outside of the sweet spots, there is an influence of states previously disregarded, specifically, $|een\rangle$ with $n$ even and $|oon\rangle$ with $n$ odd. We have calculated the microwave spectrum for some selected cases which display the effect of these states. This spectrum might be used to detect the sweet spots.
At the sweet spots of both DQDs (assumed identical), all states are doubly degenerate, due to a particular symmetry of the Hamiltonian, as explined in the Appendix.